# for preordered semifields and semirings II

### v4

@@ Abstract.
Among other applications, two companion papers use these results in order to derive limit new results in probability and information theory; one on asymptotics of random walks on topological abelian groups, and the other on the asymptotics of matrix majorization.
@@
Acknowledgements. We thank Erkka Theodor Haapasalo, Xiaosheng Mu, Tim Netzer and Péter Vrana for useful feedback, discussion and suggestions. We also thank Luciano Pomatto, Markus Schweighofer and Omer Tamuz for further discussions.
@@ 1.  Introduction
    The results of the present paper have found their first applications in probability and information theory. This comprises new formulas for the asymptotics of random walks on topological abelian groups [4] as well on the asymptotics of matrix majorization [2].
@@
In full technical detail, and using notions that will be introduced in the main text, our Vergleichsstellensatz for catalytic certificates is then the following.
Theorem 7.1.
@@
Conversely, if $ax\leq ay$ holds for some nonzero $a$, then the same spectral conditions (1.4, 1.5) are trivially implied with non-strict inequalities $\leq$ in place of $<$. The first two itemized assumptions on $S$ can be thought of as saying that although we do not require $1\geq 0$ to be the case, this should nevertheless not fail too badly, in the sense that the preorder relation $\leq$ must still be suitably large. In the final part of the theorem statement, the term semialgebra refers to the case in which $S$ comes equipped with a scalar multiplication by $\mathbb{R}_{+}$.
We do not currently have a Vergleichsstellensatz for asymptotic certificates that would apply at the same level of generality, but we do have one that applies under somewhat stronger assumptions on $S$ (which still fall short of assuming $1\geq 0$). This takes the following form, which in its conclusions precisely matches our main result of Part I.
Theorem 8.6.
@@
    Section 2 summarizes the main definitions and results of Part I, so that the present paper can be read independently of [3]. [^2] The proof of Theorem 8.6, which refers back to the proof of our Vergleichsstellensatz from Part I, is an exception to this.
@@
    Section 5 introduces multiplicatively Archimedean fully preordered semifields and analyzes their structure, resulting in a classification into five types. Remark 5.2 provides a sense in which these preordered semifields are the building blocks of all totally preordered semifields. And since the latter are a stepping stone for analyzing the structure of any preordered semifield by our first Vergleichsstellensatz from Part I (Theorem 2.1), the results of Section 5 are an important part of the structure theory of preordered semifields in general.
@@
    Section 6 introduces a certain derived preorder relation on any preordered semifield, the so-called ambient preorder. The main use of this construction is that even if $1$ and 00 are not preordered relative to each other in a given preordered semifield, they often will be with regards to the ambient preorder. With some effort, this then allows us to leverage the results of Part I, making them apply in particular to multiplicatively Archimedean fully preordered semifields as studied in Section 5. This results in two embeddings theorems, namely Proposition 6.7 and Proposition 6.8, which constitute the main results of this section.
@@
    Section 8 then proves an asymptotic Vergleichsstellensatz, namely the Theorem 8.6 quoted above. The proof is based on Theorem 7.1 and the compactness of the test spectrum (Proposition 8.5), as introduced in Definition 8.3. Example 8.7 briefly illustrates the statement in the same polynomial semiring example as before.
@@ 2.  Background from Part I
A commutative semiring is a set $S$ together with two commutative monoid structures $(S,+,0)$ and $(S,\cdot,1)$ such that the multiplication $\cdot$ distributes over the addition $+$ and $1\cdot 0=0$. Since we will not consider the noncommutative case at all, we simply use the term semiring as short for “commutative semiring”. The set of multiplicatively invertible elements in a semiring $S$ is denoted $S^{\times}$. There is a unique semiring homomorphism $\mathbb{N}\to S$, given by $n\mapsto n1$, and we often abuse notation by writing $n$ instead of $n1$. A semiring $S$ is zerosumfree if $x+y=0$ in $S$ implies $x=y=0$. It has quasi-complements if for every $x\in S$ there are $y\in S$ and $n\in\mathbb{N}$ such that $x+y=n$ (I.2.19).
@@
A preorder is a binary relation that is reflexive and transitive. We typically denote a preorder by $\leq$, where the symbols $\geq$, $<$ and $>$ have their standard induced meaning. We write $\preceq$ when another preorder symbol is needed. We also use the following two derived relations:
@@
    $\approx$ is the smallest equivalence relation contained in $\leq$. In other words, $x\approx y$ is shorthand for $(x\leq y)\>\land\>(y\leq x)$.
@@
    $\sim$ is the equivalence relation generated by $\leq$.
@@
A preordered semiring is a semiring together with a preorder relation such that for every $a\in S$, both addition by $a$ and multiplication by $a$ are monotone maps $S\to S$ (I.3.7). Note that $1\geq 0$ is not required. If $S$ is a preordered semiring, then so is $S^{\mathrm{op}}$. Preordered semifields and preordered semidomains (I.3.16/21) are preordered semirings that are semifields, respectively semidomains, with the additional condition that $x\approx 0$ implies $x=0$; for a preordered semifield, this means equivalently that $1\not\leq 0$ or $1\not\geq 0$.
@@
A preordered semifield $F$ is multiplicatively Archimedean if for all nonzero $x,y>1$ in $F$ there is $k\in\mathbb{N}$ with $y\leq x^{k}$ (Definition I.4.1). The paradigmatic examples of multiplicatively Archimedean semifields are $\mathbb{R}_{+}$ with its usual algebraic structure as well as the tropical reals $\mathbb{TR}_{+}$, defined as
@@
where the first equation is the multiplicative picture of $\mathbb{TR}_{+}$ and the second the additive picture. The isomorphism between them is given by the natural logarithm [^3] While the logarithm with respect to any base will work, we have found it convenient to fix a particular choice.. As far as it matters, we always specify explicitly which picture of $\mathbb{TR}_{+}$ we use, with the multiplicative picture often being preferred. The tropical reals also contain the Boolean semifield $\mathbb{B}=\{0,1\}$, in which $1+1=1$. It is the terminal object in the category of strict semifields and semiring homomorphisms. All the strict semifields mentioned in this paragraph are preordered semifields with respect to the standard preorder relation.
@@
Let $F$ be a multiplicatively Archimedean preordered semifield. Then $F$ order embeds into one of the following:
@@
[…]
@@
This theorem specializes to a version of Strassen’s Vergleichsstellensatz [7, Corollary 2.6] for $u=2$. In Part I, we also showed how the classical Positivstellensatz of Krivine–Kadison–Dubois can be derived from the latter, and therefore also from our Vergleichsstellensatz, in an elementary way (I.8.4).
3.  Further basic definitions
@@
A semialgebra [^5] Since we will not consider scalar multiplication by anything other than $\mathbb{R}_{+}$, we prefer omitting explicit mention of the semifield of scalars. $S$ is a semiring together with a scalar multiplication
+×$\mathbb{R}_{+}\times S$ $\longrightarrow S$
@@
In other words, a semialgebra is a semiring which at the same time is a semimodule over $\mathbb{R}_{+}$ [5, Chapter 14] in such a way that the multiplication of $S$ is bilinear. [^6] The definition of semimodule includes the additional law $r(sx)=(rs)x$, but this is implied by the ones we have assumed. As with algebras over commutative rings, a semialgebra is equivalently a semiring $S$ equipped with a semiring homomorphism $\mathbb{R}_{+}\to S$.
@@
    $x\sim y\quad\Longrightarrow\quad x\leq y\enspace\lor\enspace x\geq y$
@@
$a\leq x,y\enspace\lor\enspace x,y\leq a\quad\Longrightarrow\quad x\leq y\enspace\lor\enspace x\geq y.$
@@ 4.  Inequalities in preordered semirings and semifields
    This is trivial for $n=0$ or $m=0$. For the induction step in $n$ assuming fixed $m\geq 1$, we use
@@
$x+1\leq y+1\quad\Longrightarrow\quad x^{n}\leq y^{n+1}\enspace\forall n\in\mathbb{N}.$
@@
$\begin{matrix}a+x\leq a+y\\[2.0pt] a^{-1}+x\leq a^{-1}+y\end{matrix}\quad\Longrightarrow\quad x^{n}\leq y^{n+1}\enspace\forall n\in\mathbb{N}.$
@@ 5.  Type classification of multiplicatively Archimedean
Theorem 2.2 is an embedding theorem for multiplicatively Archimedean totally preordered semifields. We now aim at generalizing this statement to a substantially more difficult case, namely from total semifield preorders to those that are merely full in the sense of Definition 3.7, i.e. total on connected components. In particular, this does not require 00 and $1$ to be ordered relative to one another. We now explain why an embedding theorem as simple as Theorem 2.2 cannot be expected to hold.
@@
$r_{1}+s_{1}X\leq r_{2}+s_{2}X\qquad:\Longleftrightarrow\qquad r_{1}=r_{2}\enspace\land\enspace s_{1}\leq s_{2}.$
@@
5.2 Remark.
@@
$x\leq_{a}y\quad:\Longleftrightarrow\quad\mathopen{}\mathclose{{}\left(\exists k\in\mathbb{N}:x\leq ya^{k}\>\land\>y\leq xa^{k}}\right)\>\land\>\mathopen{}\mathclose{{}\left(\forall n\in\mathbb{N}:x^{n}\leq y^{n}a}\right)$
Then the layer preorder $\leq_{a}$ makes $F$ into a multiplicatively Archimedean fully preordered semifield.
@@
The $\leq_{a}$-monotonicity of multiplication is obvious, while the $\leq_{a}$-monotonicity of addition follows by $a\geq 1$ and the binomial expansion. Fullness follows by totality of $\leq$ and $x\sim_{a}y$ being equivalent to the first condition involving $k$ only. It remains to establish multiplicative Archimedeanicity. So let $x,y>_{a}1$. Then there are $k,n\in\mathbb{N}$ with $y\leq a^{k}$ and $x^{n}>a$, and hence $y\leq x^{kn}$, which produces the claim.
@@
It is clear that there can be at most one such $r$, since otherwise $x^{q}$ and $u^{p}$ would be strictly ordered in both directions for any $\frac{p}{q}$ that lies between them. Also the set of fractions $\frac{p}{q}$ with $x^{q}>u^{p}$ is easily seen to be downwards closed, and similarly the set of fractions $\frac{p}{q}$ with $x^{q}<u^{p}$ is upwards closed. Moreover, if some fraction $\frac{p}{q}$ is strictly smaller than all fractions in the first set, then it clearly must belong to the second set, and vice versa.
To see that these sets make up a Dedekind cut, it thus remains to show that they are both nonempty. But indeed $x^{q}<u^{p}$ holds for sufficiently large $\frac{p}{q}$ since $u$ is a power universal element, and similarly $x^{q}>u^{p}$ holds for small enough (sufficiently negative) $\frac{p}{q}$. ∎
@@
Assume $x>y$ for contradiction. Then there is there is $k\in\mathbb{N}$ such that $z<(xy^{-1})^{k}$ by multiplicative Archimedeanicity, and hence $y^{k}z<x^{k}\leq y^{k}z$, a contradiction. ∎
@@
For any $F$ and any nonzero $x>1$, the element $x+x^{-1}$ must lie somewhere in the order interval $[2x^{-1},2x]$. These five types thus make a distinction depending on where $x+x^{-1}$ lands in that interval, using the three elements
@@
Here, $x<y$ implies that $x^{2}+y^{2}$ must be somewhere in the order interval $[2x^{2},2y^{2}]$.
@@
    If $x\geq 1$, then $p(x)\leq q(x)$ if and only if $p(1)=q(1)$ and $p^{\prime}(1)\leq q^{\prime}(1)$.
@@
For (ii), the “if” part is immediate from (i). For the “only if” part, note first that $p(x)\sim q(x)$ is equivalent to $p(1)=q(1)$ since $p(x)\sim p(1)$ and $q(x)\sim q(1)$. Taking $x>1$ without loss of generality, assuming $p(x)\leq q(x)$ amounts to
@@
by (i). This implies $x^{p^{\prime}(1)}+1\leq x^{q^{\prime}(1)}+1$ by chaining. But then $x^{p^{\prime}(1)}\leq x^{q^{\prime}(1)}$ because of Proposition 5.10, which implies $p^{\prime}(1)\leq q^{\prime}(1)$ thanks to $x>1$.
@@
We now claim that if this holds for some $m\geq 2$, then it also holds with $m-1$ in place of $m$. Indeed the following estimates show that it is enough to increase $k$ by $1$,

### v5

@@ Abstract.
Among other applications, two companion papers use these results in order to derive new results in probability and information theory; one on asymptotics of random walks on topological abelian groups, and the other on the asymptotics of matrix majorization.
@@
Acknowledgements. We thank Erkka Theodor Haapasalo, Zhang Jiang, Milán Mosonyi, Xiaosheng Mu, Tim Netzer, Luciano Pomatto, Markus Schweighofer, Omer Tamuz, Marco Tomamichel, Becca Verghese, Frits Verhagen, Péter Vrana as well as an anonymous referee for useful feedback, discussion and suggestions.
@@ 1. Introduction
    The results of the present paper have found their first applications in probability and information theory. This comprises new formulas for the asymptotics of random walks on topological abelian groups [6] as well as for the asymptotics of matrix majorization [3].
@@
In full technical detail, and using notions that will be introduced in the main text, our Vergleichsstellensatz for catalytic certificates is then the following. We will prove it as Theorem 7.1.
1.1 Theorem.
@@
Conversely, if $ax\leq ay$ holds for some nonzero $a$, then the same spectral conditions (1.4, 1.5) are trivially implied with non-strict inequalities $\leq$ in place of <<. The first two itemized assumptions on $S$ can be thought of as saying that although we do not require $1\geq 0$ to be the case, this should nevertheless not fail too badly, in the sense that the preorder relation $\leq$ must still be suitably large. In the final part of the theorem statement, the term semialgebra refers to the case in which $S$ comes equipped with a scalar multiplication by $\mathbb{R}_{+}$.
We do not currently have a Vergleichsstellensatz for asymptotic certificates that would apply at the same level of generality, but we do have one that applies under somewhat stronger assumptions on $S$ (which still fall short of assuming $1\geq 0$). This takes the following form, which in its conclusions precisely matches our main result of Part I. It will be proven as Theorem 8.6.
1.2 Theorem.
@@
    Section 2 summarizes the main definitions and results of Part I, so that the present paper can be read independently of [5].
@@
    Section 5 introduces multiplicatively Archimedean fully preordered semifields and analyzes their structure, resulting in a classification into five types. Definition 5.3 provides a sense in which these preordered semifields are the building blocks of all totally preordered semifields. And since the latter are a stepping stone for analyzing the structure of any preordered semifield by our first Vergleichsstellensatz from Part I (Theorem 2.1), the results of Section 5 are an important part of the structure theory of preordered semifields in general.
@@
    Section 6 introduces a certain derived preorder relation on any preordered semifield, the so-called ambient preorder. The main use of this construction is that even if $1$ and 0 are not preordered relative to each other in a given preordered semifield, they often will be with regards to the ambient preorder. With some effort, this then allows us to leverage the results of Part I, making them apply in particular to multiplicatively Archimedean fully preordered semifields as studied in Section 5. This results in two embeddings theorems, namely Proposition 6.7 and Proposition 6.8, which constitute the main results of this section.
@@
    Section 8 then proves an asymptotic Vergleichsstellensatz, namely the above Theorem 1.2. The proof is based on Theorem 1.1 and the compactness of the test spectrum (Proposition 8.5), as introduced in Definition 8.3. Example 8.7 briefly illustrates the statement in the same polynomial semiring example as before.
Relation to existing literature
The lack of additive inverses in semirings frequently forces us to use very different methods than those used standardly in the proofs of Positivstellensätze. This relates to the additive structure often being considered primary in real algebra, which manifests itself e.g. in the definition of Archimedeanicity or in the use of functional-analytic methods like decompositions of states into pure states. In contrast, our methods put the emphasis on the multiplicative structure and exploit the existence of multiplicative inverses in semifields. A good case in point are the proofs of the various inequalities derived in Section 4.
Nevertheless, some of our methods overlap with those used in proofs of Positivstellensätze. This applies in particular to the following ideas:
  • $\triangleright$
    The reduction to total semifield preorders of Theorem 2.1 mirrors a very general proof technique for separation theorems, used already e.g. in Artin’s solution to Hilbert’s 17th problem [10, Lemma 1.4.4].
  • $\triangleright$
    Our use of compactness arguments for the spectrum to achieve strict is parallel to standard arguments in real algebra. Compare e.g. [10, Theorem 5.7.2] with our Section 8.
  • $\triangleright$
    Our use of derivations to capture infinitesimal information is also reminiscent of the use of derivative information, as for example in Positivstellensatz-type results of Marshall [9, Theorem 2.3], Scheiderer [11, Corollary 3.6] and Burgdorf, Scheiderer and Schweighofer [1, Theorem 7.8].
@@ 2. Background from Part I
A commutative semiring is a set $S$ together with two commutative monoid structures $(S,+,0)$ and $(S,\cdot,1)$ such that the multiplication $\cdot$ distributes over the addition ++ and $1\cdot 0=0$. Since we will not consider the noncommutative case at all, we simply use the term semiring as short for “commutative semiring”. The set of multiplicatively invertible elements in a semiring $S$ is denoted $S^{\times}$. There is a unique semiring homomorphism $\mathbb{N}\to S$, given by $n\mapsto n1$, and we often abuse notation by writing $n$ instead of $n1$. A semiring $S$ is zerosumfree if $x+y=0$ in $S$ implies $x=y=0$. It has quasi-complements if for every $x\in S$ there are $y\in S$ and $n\in\mathbb{N}$ such that $x+y=n$ (I.2.19).
@@
A preorder is a binary relation that is reflexive and transitive. We typically denote a preorder by $\leq$, where the symbols $\geq$, << and >> have their standard induced meaning. We write $\preceq$ when another preorder symbol is needed. We also use the following two derived relations:
@@
    $\approx$ is the largest equivalence relation contained in $\leq$. In other words, $x\approx y$ is shorthand for $(x\leq y)\>\land\>(y\leq x)$.
@@
    $\sim$ is the smallest equivalence relation containing $\leq$. In other words, $x\sim y$ is shorthand for the existence of a finite sequence $x=z_{0},z_{1},\ldots,z_{n}=y$ such that $z_{i}\leq z_{i+1}$ or $z_{i}\geq z_{i+1}$ for all $i$.
@@
A preordered semiring is a semiring together with a preorder relation such that for every $a\in S$, both addition by $a$ and multiplication by $a$ are monotone maps $S\to S$ (I.3.7). Note that $1\geq 0$ is not required. If $S$ is a preordered semiring, then so is $S^{\mathrm{op}}$. Preordered semifields and preordered semidomains (I.3.16/21) are preordered semirings that are semifields, respectively semidomains, with the additional condition that $x\approx 0$ implies $x=0$; for a preordered semifield, this means equivalently that $1\not\leq 0$ or $1\not\geq 0$. If $S$ is a preordered semidomain, then $\operatorname{\mathsf{Frac}}(S)$ is a preordered semifield in which
$\frac{x}{a}\leq\frac{y}{b}\quad:\Longleftrightarrow\quad\exists r\in S\setminus\{0\}:\;xb\leq ya.$ (2.1)
With this definition, $\operatorname{\mathsf{Frac}}(S)$ is a preordered semifield (I.3.22) such that the canonical homomorphism $S\to\operatorname{\mathsf{Frac}}(S)$ is monotone (I.3.23).
@@
A preordered semifield $F$ is multiplicatively Archimedean if $x^{k}\leq y$ for all $k\in\mathbb{N}$ implies $x\leq 1$ for all $x,y\in F^{\times}$ (Definition I.4.1). The paradigmatic examples of multiplicatively Archimedean semifields are $\mathbb{R}_{+}$ with its usual algebraic structure as well as the tropical reals $\mathbb{TR}_{+}$, defined as
@@
where the first equation is the multiplicative picture of $\mathbb{TR}_{+}$ and the second the additive picture. The isomorphism between them is given by the logarithm (with any base). As far as it matters, we always specify explicitly which picture of $\mathbb{TR}_{+}$ we use, with the multiplicative picture often being preferred. The tropical reals also contain the Boolean semifield $\mathbb{B}=\{0,1\}$, in which $1+1=1$. It is the terminal object in the category of strict semifields and semiring homomorphisms. All the strict semifields mentioned in this paragraph are preordered semifields with respect to the standard preorder relation.
@@
Let $F$ be a multiplicatively Archimedean totally preordered semifield. Then $F$ order embeds into one of the following:
@@
$\mathsf{TSper}(S)\coloneqq$ { $\{\textrm{ monotone hom }\phi:S\to\mathbb{R}_{+}\}$
{ $\sqcup\{\textrm{ monotone hom }\phi:S\to\mathbb{TR}_{+}\textrm{ with }\phi(u)=e\},$
@@
This theorem specializes to a version of Strassen’s Vergleichsstellensatz [12, Corollary 2.6] for $u=2$. In Part I, we also showed how the classical Positivstellensatz of Krivine–Kadison–Dubois can be derived from the latter, and therefore also from our Vergleichsstellensatz, in an elementary way (I.8.4).
3. Further basic definitions and observations
@@
3.3 Lemma.
Let $S$ be a preordered semidomain. Then the identity map induces an isomorphism of semifields
$\operatorname{\mathsf{Frac}}(S/\!\sim)\cong\operatorname{\mathsf{Frac}}(S)/\!\sim.$
Note that this slightly abuses notation by using the same symbol $\sim$ for the equivalence relation generated by the preorder on $S$ and for the one generated by the preorder on $\operatorname{\mathsf{Frac}}(S)$.
Proof.
In both semifields, the elements are given by equivalence classes of formal fractions, and we write $\equiv$ for the equivalence relation in each case. In $\operatorname{\mathsf{Frac}}(S/\!\sim)$, the equivalence relation can be described as the one generated by the following two basic equivalences:
$\frac{x}{a}\equiv\frac{rx}{ra}\>\textrm{ if }\>r\in S\setminus\{0\},\qquad\frac{x}{a}\equiv\frac{y}{a}\>\textrm{ if }\>x\leq y,\qquad\frac{x}{a}\equiv\frac{x}{b}\>\textrm{ if }\>a\leq b.$ (3.1)
On $\operatorname{\mathsf{Frac}}(S)/\!\sim$, we still have the first kind of generating equivalences, but the second and third kind is replaced by
[…]
If either of the equivalences in (3.1) holds, then it is clear that (3.2) holds as well. Conversely if (3.2) holds, then we get
$\frac{x}{a}\equiv\frac{xbr}{abr}\equiv\frac{yar}{abr}\equiv\frac{y}{b}$
with respect to (3.1). This shows that the two equivalence relations coincide, and therefore the identity map induces an isomorphism of semifields as claimed. ∎
@@
A semialgebra [^3] Since we will not consider scalar multiplication by anything other than $\mathbb{R}_{+}$, we omit explicit mention of the semifield of scalars. $S$ is a semiring together with a scalar multiplication
$\mathbb{R}_{+}\times S$ $\longrightarrow S$
@@
In other words, a semialgebra is a semiring which at the same time is a semimodule over $\mathbb{R}_{+}$ [7, Chapter 14] in such a way that the multiplication of $S$ is bilinear. [^4] The definition of semimodule includes the additional law $r(sx)=(rs)x$, but this is implied by the ones we have assumed. As with algebras over commutative rings, a semialgebra is equivalently a semiring $S$ equipped with a semiring homomorphism $\mathbb{R}_{+}\to S$.
@@
    $x\sim y\quad\Longrightarrow\quad x\leq y\kern 5.0pt\lor\kern 5.0ptx\geq y$
@@
$a\leq x,y\kern 5.0pt\lor\kern 5.0ptx,y\leq a\quad\Longrightarrow\quad x\leq y\kern 5.0pt\lor\kern 5.0ptx\geq y.$
@@
3.10 Lemma.
Let $F$ be a preordered semifield which is isomorphic to $\mathbb{Q}_{+}$ as a semifield. Then the preorder on $\mathbb{Q}_{>0}$ is one of the following:
  1. (a)
    $r\leq s$ for all $r,s\in\mathbb{Q}_{>0}$.
  2. (b)
    The usual order or its opposite.
  3. (c)
    $r\leq s$ if and only if $r=s$ for all $r,s\in\mathbb{Q}_{>0}$.
Proof.
@@ 4. Inequalities in preordered semirings and semifields
    This holds trivially for $n=0$ or $m=0$. For the induction step in $n$ assuming fixed $m\geq 1$, we use
@@
$x+1\leq y+1\quad\Longrightarrow\quad x^{n}\leq y^{n+1}\kern 5.0pt\forall n\in\mathbb{N}.$
@@
$\begin{matrix}a+x\leq a+y\\[2.0pt] a^{-1}+x\leq a^{-1}+y\end{matrix}\quad\Longrightarrow\quad x^{n}\leq y^{n+1}\kern 5.0pt\forall n\in\mathbb{N}.$
@@ 5. Type classification of multiplicatively Archimedean
Theorem 2.2 is an embedding theorem for multiplicatively Archimedean totally preordered semifields. We now aim at generalizing this statement to a substantially more difficult case, namely from total semifield preorders to those that are merely full in the sense of Definition 3.8, i.e. total on connected components. In particular, this does not require 0 and $1$ to be ordered relative to one another. We first explain why an embedding theorem as simple as Theorem 2.2 cannot be expected to hold.
@@
$r_{1}+s_{1}X\leq r_{2}+s_{2}X\qquad:\Longleftrightarrow\qquad r_{1}=r_{2}\kern 5.0pt\land\kern 5.0pts_{1}\leq s_{2}.$
@@
The following characterization of multiplicative Archimedeanicity will be slightly more convenient than the definition, and we will use it repeatedly.
5.2 Lemma.
A fully preordered semifield $F$ is multiplicatively Archimedean if and only if for all nonzero $x,y>1$ there is $k\in\mathbb{N}$ with $y\leq x^{k}$.
Proof.
Suppose that $F$ is multiplicatively Archimedean, which means that $x^{k}\leq y$ for all $k\in\mathbb{N}$ implies $x\leq 1$ for all $x,y\in F^{\times}$. Then if $x,y>1$, we in particular have $x\sim y^{k}$ for all $k\in\mathbb{N}$. Now if the claim is false, then we would have $x>y^{k}$ for all $k\in\mathbb{N}$, which by multiplicative Archimedeanicity would imply $x\leq 1$, contradicting the assumption $x>1$.
Conversely if the condition in the claim holds, let $x,y\in F^{\times}$ be given such that $x^{k}\leq y$ for all $k\in\mathbb{N}$. Then taking $k=0$ and $k=1$ gives $x\sim y\sim 1$, and hence $x\leq 1$. If $x>1$ instead of $x\leq 1$, then there would be $k\in\mathbb{N}$ with $y\leq x^{k}$ by assumption, contradicting $x^{k}\leq y$ for all $k\in\mathbb{N}$. ∎
The following derived preorder will be a useful tool for the upcoming classification theory.
5.3 Definition.
@@
$x\leq_{a}y\quad:\Longleftrightarrow\quad\mathopen{}\mathclose{{\left(\exists k\in\mathbb{N}:x\leq ya^{k}\>\land\>y\leq xa^{k}}}\right)\>\land\>\mathopen{}\mathclose{{\left(\forall n\in\mathbb{N}:x^{n}\leq y^{n}a}}\right)$ (5.1)
5.4 Lemma.
The layer preorder $\leq_{a}$ makes $F$ into a multiplicatively Archimedean fully preordered semifield.
Proof.
@@
The $\leq_{a}$-monotonicity of multiplication is obvious, while the $\leq_{a}$-monotonicity of addition follows by $a\geq 1$ and the binomial expansion. Fullness follows by totality of $\leq$ and $x\sim_{a}y$ being equivalent to the first condition involving $k$ only. It remains to establish multiplicative Archimedeanicity. So let $x,y>_{a}1$. Then there are $k,n\in\mathbb{N}$ with $y\leq a^{k}$ and $x^{n}>a$, and hence $y\leq x^{kn}$, which produces the claim. So let $x,y\in F^{\times}$ be such that $x\leq_{a}y^{k}$ for all $k\in\mathbb{N}$. ∎
5.5 Example.
Consider the semifield $F$ of linear function $r+sX$ from Example 5.1, preordered with
$r_{1}+s_{1}X\leq r_{2}+s_{2}X\qquad\Longleftrightarrow\qquad r_{1}<r_{2}\kern 5.0pt\lor\kern 5.0pt(r_{1}=r_{2}\kern 5.0pt\land\kern 5.0pts_{1}\leq s_{2}).$
This makes $F$ into a totally preordered semifield. The layer preorder $\leq_{1+X}$ is then precisely the semifield preorder from Example 5.1.
@@
It is clear that there can be at most one such $r$, since otherwise $x^{q}$ and $u^{p}$ would be strictly ordered in both directions for any $\frac{p}{q}$ that lies between them.
For any nonzero $y\sim 1$ and $t\in\mathbb{N}_{>0}$, we have $y\geq 1$ if and only if $y^{t}\geq 1$. Indeed $y\geq 1$ gives $y^{t}\geq 1$ by monotonicity of multiplication. If $y^{t}\geq 1$ instead, then $y<1$ would be the only alternative to $y\geq 1$ by fullness, and this would give $y^{t}<y^{t-1}<\ldots<y<1$, contradicting $y^{t}\geq 1$.
Looking at $x$ and $u$ now, for any $p,q\in\mathbb{Z}$ we must have exactly one of
$x^{q}>u^{p},\qquad x^{q}<u^{p},\qquad x^{q}\approx u^{p}$ (5.2)
by fullness of the preorder and $x\sim 1\sim u$. Applying the statement of the previous paragraph with $y\coloneqq x^{q}u^{-p}$ shows that which one of these cases happens is invariant under replacing $p$ and $q$ by $tp$ and $tq$ for any $t\in\mathbb{N}_{>0}$. Therefore the set of rationals
$D\coloneqq\mathopen{}\mathclose{{\left\{\frac{p}{q}\in\mathbb{Q}\>\Big|\>x^{q}>u^{p}}}\right\}$
is well-defined. Using common denominators for any two given fractions $p/q>p^{\prime}/q$ together with $u^{p}\geq u^{p^{\prime}}$ shows that $D$ is downwards closed. Similarly, the set
$U\coloneqq\mathopen{}\mathclose{{\left\{\frac{p}{q}\in\mathbb{Q}\>\Big|\>x^{q}<u^{p}}}\right\}$
is upwards closed. Moreover, if some $p/q$ is strictly smaller than all rationals in $D$, then it clearly must belong to $U$ and vice versa: again by the previous paragraph, the third alternative in (5.2) can happen for at most one equivalence class of fractions.
To see that $D$ and $U$ make up a Dedekind cut, it thus remains to show that they are both nonempty. But indeed $x^{q}<u^{p}$ holds for sufficiently large $\frac{p}{q}$ since $u$ is a power universal element, and similarly $x^{q}>u^{p}$ holds for small enough (sufficiently negative) $\frac{p}{q}$. ∎
@@
Assume $x>y$ for contradiction. Then there is $k\in\mathbb{N}$ such that $z<(xy^{-1})^{k}$ by multiplicative Archimedeanicity, and hence $y^{k}z<x^{k}\leq y^{k}z$, a contradiction. ∎
@@
For any $F$ and any nonzero $x>1$, the element $x+x^{-1}$ must lie somewhere in the order interval $2x^{-1},2x]$. These five types thus make a distinction depending on where $x+x^{-1}$ lands in that interval, using the three elements
@@
Here, $x<y$ implies that $x^{2}+y^{2}$ must be somewhere in the order interval $2x^{2},2y^{2}]$.
@@
    If $x>1$, then $p(x)\leq q(x)$ if and only if $p(1)=q(1)$ and $p^{\prime}(1)\leq q^{\prime}(1)$.
@@
For (ii), the “if” part is immediate from (i). For the “only if” part, we assume $p(x)\leq q(x)$. We first derive $p(1)=q(1)$ by showing that $p(1)\neq q(1)$ leads to a contradiction. We have
$p(1)\sim p(x)\sim q(x)\sim q(1).$
In terms of $r\coloneqq p(1)$ and $s\coloneqq q(1)$, we hence obtain $r\sim s$ although $r\neq s$. Then since $F$ is arctic, we get $\frac{r}{s}+\frac{s}{r}\approx 2$. Thus $t\coloneqq\frac{1}{2}\mathopen{}\mathclose{{\left(\frac{r}{s}+\frac{s}{r}}}\right)$ is a positive rational in $(1,\infty)$ with $t\approx 1$. Since some power $t^{n}$ exceeds $2$ in the usual order on $\mathbb{Q}_{+}$, we can write
$2=\frac{t^{n}+w}{1+w}$
for a unique $w\in\mathbb{Q}_{>0}$ and then also conclude $2\approx 1$. But then
$3x\approx 2+x^{3}\approx 1+2x^{3}\approx 3x^{2},$
which contradicts the assumed $x>1$. Therefore indeed $p(1)=q(1)$.
Now consider again $p(x)\leq q(x)$. Since $p(1)=q(1)$, we obtain
@@
by (i). This implies $x^{p^{\prime}(1)}+1\leq x^{q^{\prime}(1)}+1$ by chaining. But then $x^{p^{\prime}(1)}\leq x^{q^{\prime}(1)}$ because of Proposition 5.13, which implies the missing $p^{\prime}(1)\leq q^{\prime}(1)$ thanks to $x>1$.
@@
We now claim that if this holds for some $m\geq 2$, then it also holds with m1m-1 in place of $m$. Indeed the following estimates show that it is enough to increase $k$ by $1$,